OpenStudy (anonymous):

what is the 35th term of an arithmetic sequence where a1=-7 and a18 = 95

OpenStudy (amistre64):

using the formula for an arith seq: an = a1 + d(n-1) , and the given information we can do this. a18 = -7 +d(18-1) 95 = -7 +d(17) 102 = d(17) 102/17 = d in this case,

OpenStudy (amistre64):

a35 = -7 +(102/17)(35-1) = 197

OpenStudy (amistre64):

102/17 reduce to 6 if you wanna make it easier :)

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Wait what if the 1 is lowered down along with the 18...

OpenStudy (anonymous):

not to the power of but lowered in the equation if you know what i am talking about

OpenStudy (amistre64):

termed, yes; then a35 becomes a34 and the term that was a36 becomes a35

OpenStudy (amistre64):

so either evaluate the term at a36, or take a35 and add to it

OpenStudy (amistre64):

add 6 to it since that is the common difference between terms

OpenStudy (amistre64):

197 + 6 = 203 a36 = -7 +6(35) = 203 does that make sense?

OpenStudy (anonymous):

yes, well i think so. could you help me with one more to make sure i have it?

OpenStudy (amistre64):

i can try :)

OpenStudy (anonymous):

thanks :). okay the 27th term of an arithmetic sequence where a(termed 1) =38 and a(termed 17) =-74

OpenStudy (amistre64):

you type out the process as best you can and let me see how your doing

OpenStudy (amistre64):

start out with: an = a1 + d(n-1) ; and solve for d

OpenStudy (anonymous):

Ok wait so I used this equation right?

OpenStudy (amistre64):

or we can solve for d abstractly and then fill in the parts :) an - a1 ------ = d n-1 a17 - 38 -------- = d 17-1 -74 - 38 -------- = d 16 -112 -------- = d = -56/8 = -7 16

OpenStudy (amistre64):

its the standard equation for an arith seq, so yes

OpenStudy (anonymous):

Wait the choices i have though are -144, -20.5, -151, and -22.75

OpenStudy (amistre64):

then i suggest we go thru to the end of the solution instead of stopping halfway at -7

OpenStudy (amistre64):

d = -7 is not the same as a27

OpenStudy (amistre64):

an = a1+ d(n-1) a27 = 38 -7(27-1) = 38 -7(26) = 38 -182 = -144

OpenStudy (anonymous):

So thatcomes to be -144?

OpenStudy (amistre64):

as long as i did the math in me head right .... yes

OpenStudy (anonymous):

Thank you so much!!!!! I get most of them now. Thank you.

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks :)